OpenStudy (anonymous):

how would you find f(x) if the derivitive of f'(x)= (x+1)(x-2)(x+6)?

OpenStudy (anonymous):

\[f'(x)= x ^{3}+5x ^{2}-8x-12\] btw

OpenStudy (richyw):

integrate both sides

OpenStudy (anonymous):

how do i do that?

OpenStudy (richyw):

\[f(x)=\int x^3+5x^2−8x−12 \ dx\]

OpenStudy (richyw):

also note that \[\int f(x)+g(x)\ dx=\int f(x)dx+\int g(x)dx\]

OpenStudy (anonymous):

what does that sign mean?

OpenStudy (richyw):

what course are you taking?

OpenStudy (anonymous):

calculus and vectors. The original question states: Suppose that f is a differentiable function with derivitive f'(x). Find all the critical numbers of f, and determine whether each corresponds to a local max or min, or neither.

OpenStudy (anonymous):

You don't need f to do that. You're only interested in how f changes, i.e. in f' (and possibly the derivative of f').

OpenStudy (anonymous):

I am assuming that you only need the x-values because that question says critical numbers, not points.

OpenStudy (anonymous):

oh okay, well then thats easy